### v2

# Tailored Conversations beyond LLMs: A RL-Based Dialogue Manager

Abstract
In this work, we propose a novel framework that integrates large language models (LLMs) with an RL-based dialogue manager for open-ended dialogue with a specific goal. By leveraging hierarchical reinforcement learning to model the structured phases of dialogue and employ meta-learning to enhance adaptability across diverse user profiles, our approach enhances adaptability and efficiency, enabling the system to learn from limited data, transition fluidly between dialogue phases, and personalize responses to heterogeneous patient needs. We apply our framework to Motivational Interviews, aiming to foster behavior change, and demonstrate that the proposed dialogue manager outperforms a state-of-the-art LLM baseline in terms of reward, showing a potential benefit of conditioning LLMs to create open-ended dialogue systems with specific goals.
Tailored Conversations beyond LLMs: A RL-Based Dialogue Manager
Lucie Galland ISIR Sorbonne University Paris, France galland@isir.upmc.fr                        Catherine Pelachaud CNRS/ISIR Sorbonne University Paris, France pelachaud@isir.upmc.fr                        Florian Pecune CNRS/SANPSY Univ. Bordeaux Bordeaux, France pecune@u-bordeaux.fr
1 Introduction
In recent years, the demand for mental health services has surged, outpacing the availability of resources and creating a substantial gap in access to care Cameron et al. (2017). As a result, many patients face extended waiting periods before receiving therapy Cameron et al. (2017); Denecke et al. (2020). To address this challenge, virtual agents capable of simulating Motivational Interviewing (MI) have been proposed as a potential interim support system for individuals awaiting treatment. These agents can provide immediate assistance, particularly beneficial in therapeutic approaches requiring multiple sessions Fiske et al. (2019). However, their role is not to replace human therapists but to serve as a supplementary tool that enhances existing therapeutic interventions. Motivational Interviewing (MI) poses a particularly complex challenge for dialogue systems, traditionally addressed through intricate rule-based frameworks Prochaska et al. (2021); Olafsson et al. (2020). However, recent advances in natural language processing (NLP) have paved the way for leveraging large language models (LLMs) such as GPT-like architectures Baktash and Dawodi (2023) in such applications Steenstra et al. (2024), significantly expanding the scope of dialogue systems across various domains. Although these models exhibit remarkable language generation capabilities, they also present significant limitations, many of which can be addressed through insights from "traditional" dialogue research. In particular, LLMs often lack the controllability and structured decision-making of conventional rule-based systems, which are more predictable and interpretable Shidara et al. (2020). Rule-based domain-specific dialogue systems Hadi et al. (2024) offer notable advantages, including improved controllability, explainability, and the ability to integrate expert knowledge. However, they are typically less adaptable and more resource-intensive to develop. In contrast, LLMs demonstrate strong adaptability across domains but pose challenges in achieving control, transparency, and efficiency. Additionally, incorporating expert knowledge into LLMs often requires extensive domain-specific data Hadi et al. (2024). Notably, reinforcement learning (RL)-based dialogue managers Pecune and Marsella (2020) have shown promise in enhancing control and coherence in dialogue systems. Hence, a promising approach involves hybrid models that combine the strengths of both paradigms—leveraging the adaptability and generative capabilities of LLMs while integrating a domain-specific dialogue manager to regulate interactions Abu-Rasheed et al. (2024); Galland et al. (2024b).
In this work, we investigate a hybrid approach in which an RL-based dialogue manager governs an LLM to simulate MI dialogues, aiming to balance adaptability and control for more effective virtual therapy support.
2 Theoretical background: Motivational Interviewing (MI)
Motivational Interviewing (MI) is a therapeutic approach that emphasizes collaboration and supports behavioral change by guiding patients to explore the reasons and motivations behind their unhealthy behaviors. Complex dialogues, such as those in Motivational Interviewing (MI), evolve through distinct phases, each guided by unique long-term strategies Miller and Rollnick (2012). The dialogue usually begins with an engaging phase, where rapport is established, and patient engagement with the therapist is fostered. This is followed by a focusing phase, where core issues, their underlying causes, and the patient’s background are identified to set a clear focus for the conversation. The third phase is the evoking phase, which involves encouraging the patient’s motivation for change by eliciting and amplifying “change talk”. Finally, planning involves developing a specific, actionable plan for behavior change based on the patient’s motivation and goals. Therapists must ensure that specific objectives, such as achieving high levels of engagement, clarifying core issues, and cultivating sufficient motivation, are met before transitioning between phases. Furthermore, the process is not strictly linear, as therapists may need to revisit earlier phases depending on the patients’ evolving motivation and engagement. Individual variability in engagement and motivation necessitates a flexible approach. For a virtual therapist employing MI, effectively navigating across these phases is crucial. This requires discerning when to progress to the next phase, when to revisit earlier phases, and how to adapt the interaction to align with each patient’s unique needs and circumstances.
2.1 Patients’ Profiles in MI
Patients participating in motivational interviewing (MI) exhibit varying levels of readiness to change their behaviors. As proposed by anonymous (2024), these patients can be classified into three categories: Open-to-Change, Resistant-to-Change, and Receptive. Open-to-Change individuals demonstrate a strong willingness to modify unhealthy behaviors. Resistant-to-Change patients are generally reluctant to alter their current behaviors, showing a preference for maintaining the status quo. Receptive patients, while initially exhibiting low motivation to change, gradually develop a higher motivation to adopt healthier behaviors as the conversation progresses. These classifications capture variations in patients’ responses and therapists’ strategies, as discussed in anonymous (2024). The ability to adapt the flow of dialogue to these three patient profiles can significantly enhance the efficiency of the therapist’s dialogue model. A dialogue system for MI should be able to take into account the particular challenges that such dialogues pose. Such a dialogue system should be able to navigate across phases while being able to adapt to different profiles of users.
3 Related Work
Motivational Interviewing (MI) presents significant challenges for dialogue systems, as it necessitates both a structured progression through its four distinct phases —engagement, focusing, evoking, and planning Miller and Rollnick (2012) — and adaptability to diverse patients’ profiles anonymous (2024). While existing systems such as Woebot Prochaska et al. (2021) have demonstrated the feasibility of MI-based chatbots by incorporating therapeutic frameworks like cognitive behavioral therapy (CBT) and mindfulness, they predominantly rely on static or rule-based architectures. Steenstral et al. Steenstra et al. (2024) identified the limitations of rule-based approaches in maintaining adherence to therapeutic protocols and proposed leveraging LLMs for this application, demonstrating promising results. The advent of large language models (LLMs) has transformed dialogue generation, offering new possibilities for MI-based interactions Steenstra et al. (2024). Models such as GPT-like systems Baktash and Dawodi (2023) exhibit strong generative capabilities and adaptability across diverse applications, which can be leveraged to enhance the social dimensions of interaction. This aspect is particularly relevant, as Kanaoka et al. Kanaoka and Mutlu (2015) emphasized the critical role of social engagement and rapport in facilitating behavioral change. Recent research has underscored the importance of cognitive modeling and adaptability in dialogue systems to more accurately account for the mental states of both the agent and the user. For instance, He et al. He et al. (2024) introduced dual reasoning mechanisms that enable LLMs to incorporate contextual nuances, while Zhang et al. Zhang et al. (2020) explored interactive agent representations to improve dialogue coherence. Despite these advancements, LLMs remain constrained in terms of controllability and domain specificity Shidara et al. (2020). Recent efforts have sought to enhance control and applicability, particularly within task-oriented dialogue systems. Yao et al. Yao et al. (2023) employed reinforcement learning (RL) to optimize LLM prompting strategies; however, black-box LLMs continue to exhibit limitations in controllability and often generate repetitive responses. To address these shortcomings, Xu et al. Xu et al. (2023) demonstrated the advantages of integrating fine-tuned, smaller language models with larger LLMs, while Yu et al. Yu et al. (2023) employed Monte Carlo tree search for optimal action selection, improving both coherence and practical utility. Although these studies mark significant progress in hybrid model design, they often fall short of accommodating a broad spectrum of user profiles. Integrating the structured controllability of classical dialogue models with the generative flexibility of LLMs within an adaptive hybrid framework could facilitate more dynamic, personalized, and effective MI interactions. Reinforcement learning has been instrumental in optimizing dialogue policies and refining system behavior. Traditional RL-based approaches have primarily focused on enhancing user engagement and task success rates. Walker et al. Walker (2000) and Li et al. Li et al. (2016) demonstrated the efficacy of RL in training conversational agents for goal-oriented tasks, while Weber et al. Weber et al. (2018) illustrated its effectiveness in selecting contextually appropriate actions, such as humor or sound effects, to enrich user experiences. More advanced frameworks have incorporated both social and task-oriented rewards Pecune and Marsella (2020) or jointly trained user and dialogue policies Takanobu et al. (2020). Although these methods enable adaptation to different user profiles, they remain largely confined to task-oriented systems that rely on predefined natural language templates. In this work, we propose a novel framework that integrates LLMs with an RL-based dialogue manager to structure MI dialogues across different phases while dynamically adapting to diverse patient profiles. By synergizing structured control with generative flexibility, our approach enhances adaptability and efficiency, enabling the system to learn from limited data, transit fluidly between MI phases, and personalize responses to heterogeneous patient needs. The remainder of this paper is organized as follows: Section 4 details our proposed methodology, while Section 5 describes our evaluation environment. Section 6 presents the experimental results, and Section 7 provides an analysis and interpretation of our findings.
4 Method
This section outlines the methodology for developing a dialogue manager capable of navigating the distinct phases of Motivational Interviewing (MI) while adapting to diverse patients profile. The complete architecture is depicted in Figure 1.
4.1 Problem description
The objective of the proposed model is to predict the optimal action $a_{t}$ at each time step $t$, given the dialogue context $c$. Each action corresponds to a dialogue act representing the virtual therapist’s strategic behavior. The agent sentence is then generated by a conditioned large language model (LLM) that produces an utterance coherent with the context and realizes the selected dialogue act, as validated in Galland et al. (2024b).
Figure 1: Hierarchical architecture of the dialogue manager
4.2 Hierarchical RL for Managing Dialogue Phases
The dialogue manager employs a hierarchical reinforcement learning (HRL) framework to manage dialogue phases comprising a master policy and $N$ sub-policies. Each sub-policy is dedicated to a specific phase of the Motivational Interviewing (MI) process, while the master policy governs transitions between these phases over a fixed horizon $H$ (see Figure 1). The master policy orchestrates the dialogue by selecting the appropriate phase for the next $H$ dialogue turns based on the current master state. Meanwhile, sub-policies execute phase-specific strategies and actions, leveraging their respective sub-policy states. Notably, the master state and the sub-policy state differ in composition, as long-term planning requires distinct information from short-term decision-making. This hierarchical structure ensures dynamic and context-sensitive dialogue management, allowing real-time adjustments to both the patient’s evolving needs and the interaction context. It balances global objectives, such as increasing motivation for behavior change, with more localized goals, such as answering patient inquiries. The fixed decision interval of $H$ turns reduces the training horizon, simplifying the learning process and enhancing the model’s adaptability to diverse users by focusing on shorter-term adjustments. By leveraging hierarchical reinforcement learning (HRL), the model effectively manages the different phases of MI dialogue while minimizing adaptation complexity. The sub-policies handle local objectives, while the master policy adjusts to user-specific global goals. Once the sub-policies are trained, adapting to a new user requires fine-tuning only the master policy, which operates on a smaller horizon and action space. This approach enables efficient transitions between MI phases, ensuring that interactions remain tailored to the needs of each individual user.
4.3 Meta-Learning for User Adaptation
To facilitate rapid adaptation to new users, the master policy is trained using meta-learning techniques—specifically, the Model-Agnostic Meta-Learning (MAML) algorithm Finn et al. (2017). This approach allows the dialogue manager to generalize effectively across diverse user profiles while maintaining the flexibility to adapt quickly to novel ones. By leveraging MAML, the master policy learns shared strategies across the profiles of various patients and can fine-tune its behavior to accommodate individual patient needs, thus ensuring both robustness and personalization. In practice, after training, this meta-learned master policy serves as a well-informed initialization that approximates its behavior by averaging it across the profiles of users that have been seen. When deployed with human users for repeated interaction, it enables rapid personalization, requiring only a few interactions to adapt to the specific preferences and characteristics of a new user.
4.4 Algorithm and Training Framework
In this subsection, we present formally our algorithm and training framework.
4.4.1 Dialogue Management Algorithm
The model aims to predict the optimal dialogue act $a_{t}$ to maximize a reward function $\mathcal{R}(s_{t},a_{t})$. The system comprises a master policy $\theta$ and a set of $N$ sub-policies $\psi_{0},…,\psi_{N}$. The master policy, with a discrete action space of size $N$, determines the appropriate sub-policy to use for the next $H$ turns, while each sub-policy $\psi_{i}$ manages the dialogue acts of the corresponding phase within an action space of size $N_{da}=13$. At each time step $t$, the algorithm operates as follows. If $t\mod H=0$, the master policy selects the next sub-policy: $A_{t}=\theta(s_{t}^{master})$. Otherwise, the previous master action is reused: $A_{t}=A_{t-1}$. The sub-policy corresponding to $A_{t}$ then generates the next dialogue act: $a_{t}=\psi_{A_{t}}(s_{t})$. This action $a_{t}$ influences the environment, producing a user response and updating the state to $s_{t+1}$ and $s_{t+1}^{master}$. The detailed algorithm is presented in Algorithm 1
4.4.2 Training Framework
The training framework leverages a model-based reinforcement learning (RL) approach. A model-based approach enables efficient reuse of dialogue turns across multiple iterations as policies evolve. Specifically, we utilize the Soft Actor-Critic (SAC) algorithm Haarnoja et al. (2018), which enhances the system’s adaptability to new human users in online interactions. This approach allows for policy updates at each turn, maintaining the information from previous turns. Each training epoch targets a specific user profile and begins with cloning the master policy $\theta$. The optimization process occurs in two phases. In the first phase, the master policy $\theta$ is fixed, and the sub-policies $\psi_{0},\ldots,\psi_{n}$ are optimized using SAC. In the second phase, the sub-policies remain fixed while the cloned master policy $\theta_{\text{clone}}$ is optimized using SAC. After these optimizations, the updated policies are evaluated. Finally, the master policy $\theta$ is updated using the MAML algorithm. The complete training process is detailed in Algorithm 2.
Algorithm 1 Hierarchical Dialogue Management Algorithm
1:Input: State $s_{t}$ and $s^{master}_{t}$, Master policy $\theta$, Sub-policies $\psi_{0},\ldots,\psi_{n}$, Time horizon $H$
2:Initialize: Master policy $\theta$, Sub-policies $\psi_{0},\ldots,\psi_{n}$
3:for each time step $t$ do
4:     if $t\mod H=0$ then
5:         Compute master action $A_{t}=\theta(s^{master}_{t})$
6:     else
7:         Reuse previous master action $A_{t}=A_{t-1}$
8:     end if
9:     Select sub-policy $\psi_{A_{t}}$ based on $A_{t}$
10:     Compute dialogue act $a_{t}=\psi_{A_{t}}(s_{t})$
11:     Apply action $a_{t}$ to the environment
12:     Observe user response and new state $s_{t+1}$ and $s^{master}_{t+1}$
13:end for
Algorithm 2 Training Process for Hierarchical Dialogue Manager
1:Input: Master policy $\theta$, Sub-policies $\psi_{0},\ldots,\psi_{n}$, User simulator, Replay buffer $D_{sub}$, Replay buffer $D_{master}$
2:Initialize: Master policy $\theta$, Sub-policies $\psi_{0},\ldots,\psi_{n}$, Replay buffer $D_{sub}$, Replay buffer $D_{master}$
3:for each training epoch do
4:     Sample user profile $p$ and apply to the simulator
5:     Phase 1: Sub-policy Optimization
6:     Fix $\theta$ and for $N_{\text{sub}}$ dialogues:
7:     for each dialogue turn $t$ do
8:         Generate transition $(s_{t},a_{t},r_{t},s_{t+1},A_{t})$ and store in replay buffer $D_{sub}$
9:         Sample a batch $B$ from $D_{sub}$
10:         for each sub-policy $\psi_{i}$ do
11:              Optimize $\psi_{i}$: $\psi_{i}\leftarrow\text{SAC}(\psi_{i},B_{A_{t}=i})$
12:         end for
13:     end for
14:     Phase 2: Master Policy Optimization
15:     Clone master policy: $\theta_{\text{clone}}\leftarrow\theta$
16:     Fix $\psi_{0},\ldots,\psi_{n}$ and for $N_{\text{master}}$ dialogues:
17:     for each dialogue do
18:         Generate transition $(s_{t},a_{t},r_{t},s_{t+1})$ and store in replay buffer $D_{master}$
19:         Optimize $\theta_{\text{clone}}$: $\theta_{\text{clone}}\leftarrow\text{SAC}(\theta_{\text{clone}})$
20:     end for
21:     Evaluation
22:     Evaluate the updated policies $\theta_{\text{clone}}$ and $\psi_{0},\ldots,\psi_{n}$ on the task
23:     Master Policy Update
24:     Update the master policy $\theta$ using the MAML algorithm: $\theta\leftarrow\text{MAML}(\theta_{\text{clone}})$
25:     Empty Replay buffer $D_{sub}$ and Replay buffer $D_{master}$
26:end for
5 Evaluation Environment
In this section, we present the evaluation environment of our framework on a MI dialogue environment and compare our model with a state-of-the-art LLM baseline. We also perform ablations to demonstrate the efficiency of each of the model components, showing that incorporating knowledge on the flow of the dialogue in the dialogue manager development improves the resulting conversations.
5.1 Baseline
We use as a baseline a Nemo Instruct LLM [^1] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 prompted as validated in Steenstra et al. (2024). The prompt incorporates information on MI strategies, as well as techniques for addressing specific problems, such as Drinking, Smoking, and Sedentary Lifestyle. This approach was validated in Steenstra et al. (2024) through experiments with human participants.
5.2 Evaluation Environment
[…]
5.2.1 Simulate patients in MI
Simulating patients in MI has been explored in prior research. For instance, Galland et al. (2024a) proposed a prompt to simulate a user with a large language model (LLM) Galland et al. (2024b). This approach has been validated to produce contextually relevant, natural dialogue acts and utterances Galland et al. (2024b, a), although the differences between user profiles have not been tested and might be a limitation of this simulator. We use this user simulator to train and evaluate our dialogue manager. In the following of this article, the term user refers to this user simulator.
5.2.2 Action Space
The agent operates in a discrete action space consisting of 13 possible dialogue acts, which are categorized into task-oriented dialogue acts and socially oriented dialogue acts. Task-oriented dialogue acts include Asking for Consent or Validation, Providing Medical Education and Guidance, Planning with the Patient, Giving a Solution, Asking about Current Emotions, Inviting a Shift in Outlook, Asking for Information, and Reflection. Socially-oriented dialogue acts include Empathic reactions, Acknowledging Progress and Encouraging, Backchanneling, Greeting or Closing, and Normalizing Experiences while Providing Reassurance. This taxonomy was introduced in anonymous (2024).
5.2.3 State Space
The agent’s state space includes information from the most recent agent’s and user’s dialogue acts. User can use 9 different dialogue acts: Changing Unhealthy Behavior, Sustaining Unhealthy Behavior, Sharing Negative/Positive Feelings or Emotions, Sharing Personal Information, Realization or Understanding, Greeting or Closing, Backchanneling, and Asking for Medical Information. Additionally, the state space incorporates the current timestamp and an encoded representation of the dialogue context, which comprises the last three utterances.
5.2.4 Master State Space
The master policy’s state space is composed of an approximation of Context knowledge, Engagement approximation, and Evocation approximation. Context knowledge approximation is measured by the number of times the user employs the Sharing Personal Information dialogue act. Engagement approximation is determined by the number of times the user utilizes the Sharing Positive/Negative Feelings dialogue act. Evocation approximation is quantified by the number of uses of the Understanding or New Perspective dialogue act.
5.2.5 Reward Function
The reward function is designed to predict therapy outcomes by assigning specific values to different user dialogue acts. Prior research underscores the critical role of user responses, such as sustain talk, which is linked to poorer treatment outcomes Magill et al. (2014), and change talk, which is associated with reduced risk behaviors during follow-up assessments Magill et al. (2018). Additionally, the reward function incentivizes structured progression through the MI phases. A reward of $+5$ is assigned for Changing Unhealthy Behavior, as this represents the desired outcome, whereas a penalty of $-5$ is given for Sustaining Unhealthy Behavior, which should be discouraged. In the Engagement phase, a reward of $+50$ is granted for expressing feelings. Once at least two emotions have been expressed by the user, a reward of $+100$ is assigned for providing information in the Focusing phase. After at least two pieces of information have been shared, a reward of $+150$ is given for evoking-related dialogue acts, culminating in a reward of $+200$ for planning-related dialogue acts.
5.2.6 Episode Termination
An episode concludes after 40 turns or when the agent performs a closing action, marking the end of the dialogue interaction.
5.3 Hyperparameters
The model is trained for 41 epochs, with 455 conversations conducted per epoch. Of these, 300 are used to train the master policy, while 150 are dedicated to training the sub-policies, and 5 to evaluation. To speed up the training process, 5 conversations are performed in parallel. There are 6 sub-policies in total, and the master policy is executed every 3 turns. The user’s profile is fixed randomly at the beginning of each epoch. The sub-policies are trained with a learning rate of $10^{-4}$ and batch size of $1000$, the master policy uses a learning rate of $10^{-3}$ and batch size of $1000$, and the MAML (Model-Agnostic Meta-Learning) algorithm operates with a learning rate of $4*10^{-4}$. Each network is composed of 2 linear layers intercalated with Leaky ReLU activation functions and a hidden size of $32$. The model is trained over 16 hours using a 42GB GPU.
6 Results
Figure 2: Evolution of the reward function during training
Table 1: Experiment Results with Mean Rewards (* = Statistical difference with baseline)
Experiment Mean Reward (± SD)
(. = difference with the baseline p<.1)
Baseline 235± 106
Without MAML 233 ± 106
Without HRL 290 ± 169
Full model 394 ± 132 (.)
In this section, we present our results and ablation studies. Figure 2 illustrates the mean reward evolution across all three user profiles—Open to Change, Resistant to Change, and Hesitant—throughout the training process. At each evaluation epoch, five conversations are conducted with each user profile. Additionally, Table 1 presents the final experimental results. Our model’s reward performance surpasses that of the baseline, demonstrating that conditioning an LLM with our dialogue manager enhances the proportion of desirable dialogue acts. We observe high variability in reward values throughout the training process. This fluctuation is primarily due to the inherent stochasticity of dialogue environments, where the same agent action in a given state can lead to different user responses. Such variability introduces additional challenges to training, making consistent improvement more difficult to achieve. Moreover, due to computational constraints, five dialogues at each evaluation epoch are conducted. Such a low number cannot represent the variability of potential dialogues, thus further contributing to the observed variance. Despite this variability, an upward trend in performance is observed over the course of training. The RL model consistently outperforms the baseline in terms of average reward, demonstrating its effectiveness in promoting desirable dialogue behaviors.
6.1 Ablation Studies
We conduct two ablation studies to evaluate the impact of each design choice. In the first ablation study, we perform the same training procedure without employing MAML to train the master policy. In the second ablation study, we remove the hierarchical reinforcement learning (HRL) framework and train solely with the SAC algorithm (see Figure 2 and Table 1).
Effect of MAML
The inclusion of MAML improves the accumulated reward (see Table 1), suggesting that it enhances the learning of the master policy by explicitly accounting for variations across user profiles. Standard training can be biased by the sequence in which different user profiles are encountered, whereas MAML mitigates this by guiding the master policy toward an initialization that enables rapid adaptation to diverse users.
Effect of HRL
The effectiveness of HRL is further supported by the experimental results. Training with only the SAC algorithm leads to a lower accumulated reward, likely because the phase-based structure of MI dialogues is more challenging to capture without hierarchical modeling.
7 Interpretation
In this section, we analyze the generated dialogues and examine how our design choices influence them. We investigate the different MI phases to determine whether they emerge as expected and assess the impact of HRL. Additionally, we explore variations across user profiles and evaluate the effect of MAML on the generated dialogues.
Differences Between Phases
Figure 3: Dialogue act distribution over time, highlighting different dialogue phases for the full model. The intensity of the color is proportional to the use of the corresponding dialogue act in the turn.
To analyze the MI phases, we examine the distribution of dialogue acts across different dialogue turns. Dialogue acts associated with the Engaging phase, such as asking about emotions or sharing emotions, should be more prevalent at the beginning of the conversation, whereas those related to the Planning phase, such as providing solutions or promoting behavior change, should appear more frequently towards the end Miller and Rollnick (2012). While Engaging should occur throughout the entire dialogue, the later stages should be more focused on Planning. The phases are interwoven rather than strictly sequential Figure 3 shows that the full model employs engagement-related dialogue acts throughout the conversation and uses focusing-related actions in the last three quarters of the interaction. It also uses evoking- and planning-related action towards the end of the interaction. However, the HRL ablation only uses a singular action: "Give solution". This strategy may provide immediate rewards with the simulated users, but does not adhere to the MI principle, making the full model with HRL more diverse and MI compliant.
Differences between user profiles
Figure 4: Distribution of phase activation over time across different user profiles. The intensity of the color is proportional to the use of the phase in the turn.
In this section, we examine the impact of meta-learning on training the master policy. Figure 4 illustrates the distribution of master actions activation over time for different user profiles for the full model. For the model trained without meta-learning, we observe a collapse of the master policy to a single dominant action across interactions. This highlights the difficulty of learning a generalized policy that performs well across diverse user profiles without explicit mechanisms for adaptation. In contrast, the use of meta-learning allows the master policy to maintain variability and adaptability in its actions. In Figure 4, we associate each master action with MI phases based on the predominant dialogue acts typically observed in those phases and master actions, providing interpretability into the policy’s learned structure. The models effectively differentiate between distinct phases, initially engaging in Engaging/Focusing phases (Master action 5) before transitioning into Focusing/Planning phases (Master action 3) and Evoking phase (Master action 4).The engaging phase is quicker for Open to change users, as they do human therapist, aligning with prior findings anonymous (2024). Indeed, open-to-change patients require less rapport-building with the therapist, as they choose to initiate the conversation. This qualitative analysis highlights the benefits of incorporating meta-learning, as it enhances the model’s ability to structure MI phases effectively and provide a more personalized dialogue experience.
8 Conclusion
In this paper, we present a dialogue manager for MI dialogue design, addressing the unique challenges posed by this profile of interaction. We leverage HRL to model the structured phases of MI and employ meta-learning to enhance adaptability across diverse user profiles. Our findings demonstrate that the proposed dialogue manager outperforms an LLM baseline in terms of reward. Additionally, our analysis of the generated conversations provides valuable insights into how HRL and meta-learning contribute to the structured yet adaptive nature of the dialogue.
9 Limitations
The current framework is trained using a single implementation of a simulated user, which limits its generalizability. To fully assess its effectiveness, the model is currently being tested with human participants to capture a broader range of user behaviors and characteristics.
Moreover, the simulator is based on a Mistral LLM, which requires significant processing time to generate user behaviors. This limitation constrains the training capacity, as complex reinforcement learning (RL) problems like this one require extensive trial and error. As a result, certain design choices—such as small batch sizes and a limited number of conversations per epoch—were necessary, contributing to the observed training instability. Addressing this issue in future work could lead to more robust and efficient training.
Additionally, the analysis of dialogue phases currently relies on predefined heuristics, making it inherently subjective. A more rigorous approach would involve annotation and validation by professional MI annotators to ensure alignment with clinical practices, thereby improving the system’s reliability.
10 Ethical Implications
This work introduces a dialogue manager designed for Motivational Interviewing interactions. Its objective is not to replace therapists but to provide supplementary support or serve as an introduction to therapy. The focus of this research is exclusively on the dialogue management component, which operates within a constrained set of possible actions. All the LLMs are run locally, and no sensitive information is sent to outside services.
Given the sensitive nature of such applications, careful examination and validation of the language model outputs remain essential. It is imperative to emphasize in both the codebase and accompanying documentation that these interactions are not intended to replace professional therapists but rather to complement their efforts in appropriate contexts. Ensuring transparency and adherence to ethical standards is fundamental to responsibly deploying this technology. Additionally, it would be important to validate the approach with MI practitioners before using it in any real world context.
11 Acknowledgement
This work was partially funded by the ANR-DFG-JST Panorama and ANR-JST-CREST TAPAS (19-JSTS-0001-01) projects